# Generating Infinitely Many Hyperbolic Knots with Plats

### v1

@@ Abstract.
In this paper we study the relationships between links in plat position, the dynamics of the braid group, and Heegaard splittings of double branched covers of $S^{3}$ over a link. These relationships offer new ways to view links in plat position and a new tool kit for analyzing links. In particular, we show that the Hempel distance of the Heegaard splitting of the double branched cover obtained from a plat is a lower bound for the Hempel distance of that plat. Using the Hempel distance of a knot in bridge position and pseudo-Anosov braids we obtain our main result: a construction of infinitely many sequences of prime hyperbolic $n$-bridge knots for $n\geq 3$, infinitely many of which are distinct. We consider known results to show that the knot genus and hyperbolic volume of these knots are bounded below by a linear function. As a further result we show that the plat closure of a 6-braid is generically hyperbolic.
@@ 1. Introduction
Until recently, it was widely believed that generic prime links are hyperbolic. Indeed, this holds for prime knots of small crossing number [16]. However, Malyutin proved in [19] that, counter-intuitively, the proportion of hyperbolic links as crossing number increases does not approach 1. This leads to a question: how can one construct prime hyperbolic knots of high crossing number? This question has been approached, for instance, in [2] by modifying hyperbolic links with at least $2$ components combinatorially.
In this paper, we approach this question by studying knots as plat closures of elements of the braid group (called plats). By Schubert’s work in [25], every link in $S^{3}$ admits a bridge position for some $n$, and therefore can be represented as the plat closure of a word in the braid group $B_{2n}$. Links in plat position admit a measure called the Hempel distance, which is an analog of the Hempel distance for Heegaard splittings of $3$–manifolds [14]. Bachman and Schliemer showed in [10] that if the Hempel distance of a knot in bridge position is larger than $3$, then the knot is hyperbolic. This gives us a criteria to look for when constructing hyperbolic knots.
@@
Proposition 1.1.
If $(K,S)$ is a bridge splitting of a knot in $S^{3}$ and $H=H_{+}\cup_{\phi}H_{-}$ is the Heegaard splitting for the double cover of $S^{3}$ over $K$, $Y(K)$. Then $d(Y,H)\leq d(K,S)$.
@@
Using Proposition 1.1, we can then construct a family of $n$–bridge hyperbolic knots whose Hempel distance is bounded below by a linear function in $m.$ These knots are certain powers of a fixed pseudo-Anosov braid $\beta.$ If $\beta$ is chosen to have only positive crossings, the crossing numbers of this family of hyperbolic knots also increases. This provides a method for generating high crossing number hyperbolic knots. Of particular interest, our family of knots have unbounded volume. It has been shown by Bachman and Schleimer [10], that the genus of a knot is bounded below by a linear function in the Hempel distance of the knot, see Corollary 2.6. The genus of a hyperbolic knot is known to be linearly related to the volume of the knot [9]. Thus, the volume of a hyperbolic knot bounded below by a linear function in the the Hempel distance. A key feature of our sequence of plat closures is that the Hempel distance increases with each power, so that the volume of the hyperbolic knots increases as well.
Theorem 1.2.
@@
For $n\geq 3,$ if $\beta\in B_{2n}$ is a generic, pseudo-Anosov braid with $1$-component plat closure, then plat closures of certain powers of $\beta$ will generate a sequence of infinitely many distinct prime hyperbolic knots whose volumes are strictly increasing.
Birman and Hilden showed that the mapping class group of the genus 2 surface is isomorphic to the braid group $B_{6}$ [6]. Using known results about the Hempel distance of Heegaard splittings and the work of this paper, we show:
Theorem 1.3.
The plat closures of braids in $B_{6}$ are generically hyperbolic.
That is, while the conjecture that prime knots are generically hyperbolic is false, for prime knots which can be represented as the plat closure of a braid in $B_{6}$ the conjecture holds.
@@
The paper is outlined as follows: In Section 2 we introduce links in plat position and discuss the Nielsen-Thurston classification of braids. In particular, we consider pseudo-Anosov braids. Then we discuss the Hempel distance for both Heegaard splittings and links in bridge position and recall some facts about how the distance of a link in bridge position relates to its link type. In Section 3 we recall the method for lifting a link in plat position to obtain a 3-manifold and a Heegaard splitting that corresponds to the link. We show how to “see” the gluing map of the lift by considering a shadow diagram of the plat. We then prove our main proposition, the Hempel distance of a link is bounded below by the Hempel distance of the splitting map of the lift. In Section 4, prove our main theorem. We show an explicit sequence that generates infinitely many hyperbolic knots whose genus and hyperbolic volume are bounded below by linear functions. In Section 5, we show that plat closures of 6-braids are generically hyperbolic. Our final section provides some remaining questions and a few ideas for further research.
@@ Acknowledgements
We thank William Menasco and Çağatay Kutluhan for their constant guidance and suggestions on this research, as well as Deepisha Solanki, Greg Vinal, and Román Aranda for taking an active interest in this research and for many insightful conversations. We would also like to thank Dan Margalit and Saul Schleimer for their suggestions on early drafts of this manuscript. Carolyn Engelhardt was supported in part by a Simons Foundation grant No. 519352.
@@ 2.1.3. Pseudo-Anosov Braids
    Reducible: there exists a collection of disjoint curves on the punctured sphere, fixed by $\beta.$
@@
In this work, we are primarily interested in the plat closures of pseudo-Anosov braids. By Theorem 5.1 of [11], pseudo-Anosov braids are generic in $B_{n}$ in the following sense: if we consider all braid words in $B_{n}$ of length at most $\ell$, then the proportion of pseudo-Anosov braids in this collection increases exponentially as $\ell$ increases.
Psuedo-Anosov braids have many interesting properties; in particular, they leave invariant a foliation on the punctured disk. The unstable foliation in turn gives rise to an invariant train track, which is an embedded 1-manifold. See Section 3.3 for further discussion and Example 4.2 for an example of a train track on a disk. Also of note: pseudo-Anosov braids admit a numerical measure called entropy which is the logarithm of the dilation $\lambda$. Heuristically, one can think of braids as ways to “mix up” the disk, and pseudo-Anosov braids as ways to “mix up” the disk as much as possible. The entropy of a pseudo-Anosov braid can then be thought of as measuring how quickly the disk is mixing.
@@ 2.3. Hempel Distance of Links
Given a link $L$ in bridge position with respect to a bridge sphere $S$, take a point of $L\cap S$ and perturb a neighborhood of the point slightly to add two new critical points to $L$ (a maximum and a minimum). The resulting link $L^{\prime}$ is now in $(n+1)$-bridge position because $S$ is still a bridge sphere for $L^{\prime}$ and $L^{\prime}$ now intersects $S$ in $2(n+1)$ points. This move is called a stabilization of $L$. On the other hand, cancelling a maximum and a minimum by an isotopy in a neighborhood of a point of $L$ intersecting the bridge sphere $S$ is called a destabilization. An example is illustrated in Figure 6. If $L$ can be destabilized we say that it admits a destabilization.
@@ 3. Plats and the Double Branched Cover
In this section, we utilize the fact that a link $L$ in plat position is well suited to explicitly construct the double branched cover of $S^{3}$ over $L$ with a Heegaard splitting inherited from $L$. We then relate the Hempel distance of $(Y,H)$ and of $(S^{3},L)$, proving Proposition 3.5 and Lemma 3.9, which we will use in our construction in Section 4.
@@ 3.1. Constructing the Double Branched Cover
A shadow diagram on $X_{2n}$ is a collection $\mathcal{S}=\{a_{1},...,a_{n}\}$ of properly embedded pairwise disjoint arcs, each of which is called a shadow. Given any trivial tangle $T$, there is a shadow diagram such that $T$ is isotopic on to $\mathcal{S}$; in this case, we say that $\mathcal{S}$ is a shadow diagram fro $T$. In general, there are many such shadow diagrams; see a forthcoming paper of the first author for further discussion. If $\mathcal{S}=\{a_{i}\}_{i=1}^{n}$ is a shadow diagram for $T=\{t_{i}\}_{i=1}^{n}$ such that $t_{i}$ is isotopic onto $a_{i}$ for $1\leq i\leq n$, then each pair $t_{i},a_{i}$ cobound an embedded disk in $B_{+}$ called a bridge disk. We fix a shadow diagram $\mathcal{S}_{\varepsilon}$ for $\overline{T}_{\varepsilon}$ as in Figure 7. We choose any shadow diagram $\mathcal{S}_{\beta}$ which $T_{\beta}$ is isotopic to.
@@ 3.2. The Double Branched Cover and Hempel Distance
These results lead us to the following proposition:
Proposition 3.5.
If $(K,S)$ is a bridge splitting of a knot in $S^{3}$ and $H=H_{+}\cup_{\phi}H_{-}$ is the Heegaard splitting for the double cover of $S^{3}$ over $K$, $Y(K)$. Then $d(Y,H)\leq d(K,S)$.
@@ 3.3. Detecting Generic Pseudo-Anosov Braids via Train-Tracks
Given a closed surface $F$ of genus $g\geq 2$. In order to construct the knots in Theorem 4.1, we will rely on pseudo-Anosov maps $\phi\colon F\rightarrow F$ such that, as splitting maps, the Heegaard splittings determined by powers $\phi^{m}$ have increasing Hempel distance as $m$ increases. However, not all pseudo-Anosov maps have this property. In fact, there exist pseudo-Anosov maps that fix a collection of meridian curves in the handlebody [21]. An explicit example of such an element is given in [8].
@@
This is a direct consequence of Proposition 3.5. ∎
@@ 4. Generating a Sequence of Hyperbolic Knots
Theorem 4.1.
For $n\geq 3,$ if $\beta\in B_{2n}$ is a generic, pseudo-Anosov braid with $1$-component plat closure, then plat closures of certain powers of $\beta$ will generate a sequence of infinitely many distinct prime hyperbolic knots.
@@
Because $\beta$ is a generic pseudo-Anosov braid, by Lemma 3.8 and Lemma 3.9 the lift of $\beta$ will be a generic pseudo-Anosov map of the surface $F_{n-1}$. Thus, there is some $M$ such that for all $m\geq M,$ the Hempel distance of the Heegaard splitting obtained from the lifting $\beta^{mk+1}$ is larger than 3 by Lemma 3.10. By Proposition 3.5, $d(\widehat{\beta^{mk+1}},S)\geq 3.$ By Theorem 2.3, $\widehat{\beta^{mk+1}}$ is hyperbolic. We also know from Section 2.1.2 that $\widehat{\beta^{mk+1}}$ is a knot for all $m$. By Theorem 2.4, each $\widehat{\beta^{mk+1}}$ is a prime knot.
@@
We now present an example of a family of high-crossing hyperbolic knots as in Theorem 4.1.
@@
Therefore, taking the plat closure of powers of $\beta$ and applying Theorem 4.1 generates an infinite sequence of prime hyperbolic knots, infinitely many of which are distinct.
Below we provide a table showing the relationship between powers of $\beta$, the entropy, the knot genus, and the hyperbolic volume. This data was calculated in Sage and SnapPy.

### v2

@@ Abstract.
In this paper we study the relationships between links in plat position, the dynamics of the braid group, and Heegaard splittings of double branched covers of $S^{3}$ over a link. These relationships offer new ways to view links in plat position and a new tool kit for analyzing links. In particular, we show that the Hempel distance of the Heegaard splitting of the double branched cover obtained from a plat is a lower bound for the Hempel distance of that plat. Using the Hempel distance of a knot in bridge position and pseudo-Anosov braids we obtain our main result: a construction of infinitely many sequences of prime hyperbolic $n$-bridge knots for $n\geq 3$, infinitely many of which are distinct. We consider known results to show that the knot genus and hyperbolic volume of these knots are bounded below by a linear function.
@@ 1. Introduction
Until recently, it was widely believed that generic prime links are hyperbolic. Indeed, this holds for prime knots of small crossing number [16]. However, Malyutin proved in [20] that, counter-intuitively, the proportion of hyperbolic links as crossing number increases does not approach 1. This leads to the question: How can one construct prime hyperbolic knots of high crossing number? This question has been approached, for instance, in [2] by modifying hyperbolic links with at least $2$ components combinatorially.
In this paper, we approach this question by studying knots as plat closures of elements of the braid group (called plats). By Schubert’s work in [27], every link in $S^{3}$ admits a bridge position for some $n$, and therefore can be represented as the plat closure of a word in the braid group $B_{2n}$. Links in plat position admit a measure called the Hempel distance, which is an analog of the Hempel distance for Heegaard splittings of $3$–manifolds [14]. Bachman and Schliemer showed in [10] that if the Hempel distance of a knot in bridge position is greater than $3$, then the knot is hyperbolic. This gives us a criterion to look for when constructing hyperbolic knots.
@@
{prop}
[] If $(K,S)$ is a bridge splitting of a knot in $S^{3}$ and $H=H_{+}\cup_{\phi}H_{-}$ is the Heegaard splitting for the double cover of $S^{3}$ over $K$, $Y(K)$. Then $d(Y,H)\leq d(K,S)$.
@@
Using Proposition 1, we can then construct a family of $n$–bridge hyperbolic knots whose Hempel distance is bounded below by a linear function in $m$ by fixing a pseudo-Anosov braid word $\beta$ which meets certain technical conditions and taking the plat closures of $\beta^{m}$.
{thm}
@@
[] For $n\geq 3,$ if $\beta\in B_{2n}$ is a generic, pseudo-Anosov braid with $1$-component plat closure, then plat closures of certain powers of $\beta$ will generate a sequence of infinitely many distinct prime hyperbolic knots whose volumes are strictly increasing.
In particular, if $\beta$ is chosen to have only positive crossings, the crossing numbers of this family of hyperbolic knots also increases. This provides a method for generating high crossing number hyperbolic knots. Notably, the Hempel distance of $k$ in such a family is bounded below by a linear function in $m$, but we do not have an upper bound for the Hempel distance. In contrast, in [17] Moriah and Johnson calculated precisely the Hempel distance for plat closures of braid words of a certain form, called highly-twisted knots. Our construction does not require that the plat projections are highly-twisted, as in Examples 4.1 and 4.5 (though we do not know if these admit highly-twisted diagrams in the sense of [19]). That is, our construction is slightly more general, but loses the sharpness of [17]. However, there is clear synergy; in Example 4.4, we construct a pseudo-Anosov braid whose plat closure is a highly-twisted knot and use the results of Moriah and Johnson to explicitly calculate the Hempel distances of the plat closures of powers of this braid word.
Also of interest, our family of knots have unbounded volume. It has been shown by Bachman and Schleimer in [10] that the genus of a knot is bounded below by a linear function in the Hempel distance of the knot (see Corollary 2.6). The genus of a hyperbolic knot is known to be linearly related to the volume of the knot [9]. Thus, the volume of a hyperbolic knot is bounded below by a linear function in the Hempel distance. A key feature of our sequence of plat closures is that the Hempel distance increases with each power, so that the volume of the hyperbolic knots increases as well.
@@
The paper is outlined as follows: in Section 2 we introduce links in plat position and discuss the Nielsen-Thurston classification of braids. In particular, we consider pseudo-Anosov braids. Then we discuss the Hempel distance for both Heegaard splittings and links in bridge position and recall some facts about how the distance of a link in bridge position relates to its link type. In Section 3 we recall the method for lifting a link in plat position to obtain a 3-manifold and a Heegaard splitting that corresponds to the link. We show how to “see” the gluing map of the lift by considering a shadow diagram of the plat. We then prove our main proposition, the Hempel distance of a link is bounded below by the Hempel distance of the splitting map of the lift. In Section 4, prove our main theorem. We show an explicit sequence that generates infinitely many hyperbolic knots whose genus and hyperbolic volume are bounded below by linear functions. Our final section provides some remaining questions and a few ideas for further research.
@@ Acknowledgements
We thank William Menasco and Çağatay Kutluhan for their constant guidance and suggestions on this research, as well as Deepisha Solanki, Greg Vinal, and Román Aranda for taking an active interest in this research and for many insightful conversations. We would also like to thank Dan Margalit and Saul Schleimer for their suggestions on early drafts of this manuscript. We thank for Ryan Blair, Nathan Dunfield, Nir Lazarovich, Eric Samperton, and Moirah Yoav for their helpful correspondence and insights into the first draft of this paper. Carolyn Engelhardt was supported in part by a Simons Foundation grant No. 519352.
@@ 2.1.3. Pseudo-Anosov Braids
    Reducible: there exists a collection of disjoint curves on the punctured sphere fixed by $\beta.$
@@
In this work, we are primarily interested in the plat closures of pseudo-Anosov braids. By Theorem 5.1 of [11], pseudo-Anosov braids are generic in $B_{n}$ in the following sense: if we consider all braid words in $B_{n}$ of length at most $\ell$, the proportion of pseudo-Anosov braids in this collection increases exponentially as $\ell$ increases.
Pseudo-Anosov braids have many interesting properties; in particular, they leave invariant a foliation on the punctured disk. The unstable foliation in turn gives rise to an invariant train track, which is an immersed 1-manifold. See Section 3.3 for further discussion and Example 4.1 for an example of a train track on a disk. Also of note: pseudo-Anosov braids admit a numerical measure called entropy which is the logarithm of the dilation $\lambda$. Heuristically, one can think of braids as ways to “mix up” the disk, and pseudo-Anosov braids as ways to “mix up” the disk as much as possible. The entropy of a pseudo-Anosov braid can then be thought of as measuring how quickly the disk is mixing.
@@ 2.3. Hempel Distance of Links
Given a link $L$ in bridge position with respect to a bridge sphere $S$, take a point of $L\cap S$ and perturb a neighborhood of the point slightly to add two new critical points to $L$ (a maximum and a minimum). The resulting link $L^{\prime}$ is now in $(n+1)$-bridge position because $S$ is still a bridge sphere for $L^{\prime}$ and $L^{\prime}$ now intersects $S$ in $2(n+1)$ points. This move is called a stabilization of $L$. On the other hand, canceling a maximum and a minimum by an isotopy in a neighborhood of a point of $L$ intersecting the bridge sphere $S$ is called a destabilization. An example is illustrated in Figure 6. If $L$ can be destabilized we say that it admits a destabilization.
@@ 3. Plats and the Double Branched Cover
In this section, we utilize the fact that a link $L$ in plat position is well suited to explicitly construct the double branched cover of $S^{3}$ over $L$ with a Heegaard splitting inherited from $L$. We then relate the Hempel distance of $(Y,H)$ and of $(S^{3},L)$, proving Proposition 1 and Lemma 3.9, which we will use in our construction in Section 4.
@@ 3.1. Constructing the Double Branched Cover
A shadow diagram on $X_{2n}$ is a collection $\mathcal{S}=\{a_{1},...,a_{n}\}$ of properly embedded pairwise disjoint arcs, each of which is called a shadow. Given any trivial tangle $T$, there is a shadow diagram such that $T$ is isotopic on to $\mathcal{S}$; in this case, we say that $\mathcal{S}$ is a shadow diagram fro $T$. In general, there are many such shadow diagrams; see a forthcoming paper of the first author for further discussion. If $\mathcal{S}=\{a_{i}\}_{i=1}^{n}$ is a shadow diagram for $T=\{t_{i}\}_{i=1}^{n}$ such that $t_{i}$ is isotopic onto $a_{i}$ for $1\leq i\leq n$, then each pair $t_{i},a_{i}$ cobound an embedded disk in $B_{+}$ called a bridge disk. We fix a shadow diagram $\mathcal{S}_{\varepsilon}$ for $\overline{T}_{\varepsilon}$ as in Figure 7. We choose any shadow diagram $\mathcal{S}_{\beta}$ which $T_{\beta}$ is isotopic onto.
@@ 3.2. The Double Branched Cover and Hempel Distance
Remark 3.4.
We note that while $\mathcal{AC}(X_{2n})$ embeds into $\mathcal{C}(F_{n-1})$, they are not isomorphic. Consider a separating curve on $F_{2}$; this projects to an inessential curve on $X_{2n}$.
@@
These results lead us to our main proposition:
See 1
@@ 3.3. Detecting Generic Pseudo-Anosov Braids via Train-Tracks
Given a closed surface $F$ of genus $g\geq 2$. In order to construct the knots in Theorem 1, we will rely on pseudo-Anosov maps $\phi\colon F\rightarrow F$ such that, as splitting maps, the Heegaard splittings determined by powers $\phi^{m}$ have increasing Hempel distance as $m$ increases. However, not all pseudo-Anosov maps have this property. In fact, there exist pseudo-Anosov maps that fix a collection of meridian curves in the handlebody [22]. An explicit example of such an element is given in [8].
@@
This is a direct consequence of Proposition 1. ∎
@@ 4. Generating a Sequence of Hyperbolic Knots
See 1
@@
Because $\beta$ is a generic pseudo-Anosov braid, by Lemma 3.8 and Lemma 3.9 the lift of $\beta$ will be a generic pseudo-Anosov map of the surface $F_{n-1}$. Thus, there is some $M$ such that for all $m\geq M,$ the Hempel distance of the Heegaard splitting obtained from the lifting $\beta^{mk+1}$ is larger than 3 by Lemma 3.10. By Proposition 1, $d(\widehat{\beta^{mk+1}},S)\geq 3.$ By Theorem 2.3, $\widehat{\beta^{mk+1}}$ is hyperbolic. We also know from Section 2.1.2 that $\widehat{\beta^{mk+1}}$ is a knot for all $m$. By Theorem 2.4, each $\widehat{\beta^{mk+1}}$ is a prime knot.
@@
Therefore, taking the plat closure of powers of $\beta$ and applying Theorem 1 generates an infinite sequence of prime hyperbolic knots, infinitely many of which are distinct. In Table 1 we provide a table showing the relationship between powers of $\beta$, the entropy, the knot genus, and the hyperbolic volume. This data was calculated in Sage and SnapPy.
Table 1. The entropy, genus, and volume for $\widehat{\beta^{n}}$ when $\beta=\sigma_{2}^{2}\sigma_{4}\sigma_{1}\sigma_{3}\sigma_{5}\sigma_{2}$
@@
4.1. Relationship to Highly Twisted Links
In [17], Johnson and Moriah consider consider plats which have the form shown in Figure 13 below, which we will refer to as fishnet plats. Each of the boxes is called a twist box and is denoted by $a_{i,j}$ for $11\leq i\leq n-$ 1 and $1\leq j\leq m.$ The integer $n$ is referred to as the height of the plat, and $m$ is called the width (in our work, $m$ is the bridge number of the plat). Each twist box is assigned an integer that denotes how many half twists occur in the box, and whether they are right-handed (positive) or left-handed (negative). Such a diagram is highly twisted if $|a_{i,j}|\geq 3$ for every $a_{i,j}$.
Figure 13. On the left is a general diagram for a height 6 width 3 fishnet plat, as described in [17]. On the right is an example of a highly twisted plat with $a_{1,j}=a_{3,j}=3$ for $j=1,2$ and $a_{2,j}=-3$ for $j=1,2,3$.
While every link admits a presentation as a fishnet plat, not every link admits a presentation as a highly twisted plat. For highly twisted plats, Johnson and Moriah prove the following:
Theorem 4.2 (Theorem 1.2 of [17]).
If $K\subseteq S^{3}$ is an $m$–bridge link with a highly twisted $n$–row $2m$–plat projection, with $m\geq 3,$ then the distance of the Hempel distance of the plat is exactly $\left\lceil\frac{n}{2(m-2)}\right\rceil.$
In [19], Lazarovich, Moriah, and Pinksy expand upon this result to knots and links which are not necessarily in plat position, showing:
Theorem 4.3 (Theorem A of [19]).
Let $D(\mathcal{L})$ be a connected, prime, twist-reduced, $3$-highly twisted diagram of a link $\mathcal{L}$ with at least two twist regions, then $\mathcal{L}$ is hyperbolic.
There is a deep connection between hyperbolic knots and knots which admit a highly-twisted diagram. However, when in plat position, the restriction to highly-twisted braid words is very strict; braid words which are highly twisted are of a very specific form. In our Example 4.1, $\beta$ is not a highly-twisted diagram, nor are powers of $\beta$. However, this does not preclude the possibility that these knots do admit some highly-twisted diagram.
Example 4.4.
While our construction does not require highly-twisted braid words, it does interact very nicely with them. For example, we consider the braid word: $\beta=\sigma_{2}^{3}\sigma_{4}^{3}\sigma_{1}^{-3}\sigma_{3}^{-3}\sigma_{5}^{-3}\sigma_{2}^{3}\sigma_{4}^{3}$ as in Figure 13. First, we show this braid word is pseudo-Anosov by considering its lift to $Y(\widehat{\beta})$. Lifting $\sigma_{2}$ and $\sigma_{4}$ to their corresponding action on the genus 2 splitting surface for $Y(\widehat{\beta})$, we see that they correspond to a multi-twist along the multicurve containing the two longitudinal curves. Similarly, the letters $\sigma_{1},\sigma_{3},\sigma_{5}$ lift to a multi-twist along the multicurve containing the two meridional curves and a curve that intersects each longitude exactly once. This is illustrated in Figure 14. Since the union of the two multicurves fill the surface, it follows from work of Penner [26] that taking positive powers of twists along one multicurve and negative powers of twists along the other multicurve results in a pseudo-Anosov map on the genus 2 surface. Subsequently $\beta$ is a pseudo-Anosov map on the punctured sphere since $\beta$ and its lift have the same local properties, i.e. the same stable and unstable foliations in a small neighborhood. (For more on Penner’s construction, see Theorem 14.4 of [12].) Therefore, $\beta$ is a pseudo-Anosov braid word.
Further, we see that $\beta$ must be generic in the sense of Section 3.3 because its plat closure is highly twisted; all subsequent powers of $\beta$ will result in highly twisted plats, and by Theorem 4.2 the Hempel distance of these powers must then increase linearly.
Figure 14. The multicurve corresponding to lifts of $\sigma_{2}$ and $\sigma_{4}$ in blue. The multicurve corresponding to lifts of $\sigma_{1},\sigma_{3},\sigma_{5}$ in red.
We chose our braid $\beta$ specifically so that it would be both highly twisted and fit the hypotheses of Theorem 1. The associated permutation of $\beta$ is $\pi(\beta)=[3,5,1,6,2,4].$ We see by constructing the plat closure graph that $\beta$ is a knot. The order of the permutation $\pi(\beta)$ is 2. Therefore, $\{\widehat{\beta^{2k+1}}|k=0,1,2,\dots\}$ is a collection of knots. Using Theorem 4.2 we can calculate explicitly the Hempel distances of these plats:
$d(\widehat{\beta^{2k+1}})$ $=\left\lceil\frac{n}{(2(m-2)}\right\rceil\text{\hskip 10.00002pt for all $k,$ $m=3$ }$
$=\left\lceil\frac{n}{2}\right\rceil\text{\hskip 10.00002pt for each $k,$ $n=4(k+1)$ }$
$=\left\lceil\frac{4(k+1)}{2}\right\rceil=2(k+1)$
This demonstrates that the Hempel distance grows linearly with powers of $\beta$ as stated in Lemma 3.10. In Table 2 we provide a table showing the first few knotted powers of $\beta,$ the entropy, the Hempel distance, and hyperbolic volume. Note that we were unable to calculate the knot genus for powers of this knot due to computational limitations. This data was calculated in Sage and SnapPy.
Table 2. The entropy, genus, and volume for $\widehat{\beta^{n}}$ when $\beta=\sigma_{2}^{3}\sigma_{4}^{3}\sigma_{1}^{-3}\sigma_{3}^{-3}\sigma_{5}^{-3}\sigma_{2}^{3}\sigma_{4}^{3}$
$k$ Entropy Hempel distance Hyperbolic volume
0 4.02503 2 28.75
1 12.07510 4 90.49
2 20.12515 6 152.20
3 28.17521 8 213.92
4 36.2257 10 275.63
5 44.27533 12 337.35
6 52.32539 14 399.06
7 60.37545 16 460.78
8 68.42551 18 522.49
Example 4.5.
By the same argument as in Example 4.4, the braid word $\beta=\sigma_{2}^{3}\sigma_{4}^{3}\sigma_{1}^{-1}\sigma_{3}^{-1}\sigma_{5}^{-1}\sigma_{2}^{3}\sigma_{4}^{3}$ is a pseudo-Anosov braid. By lifting its train-track, we see that it is generic in the sense of Section 3.3. The permutation $\pi_{\beta}=[3,5,1,6,2,4]$ is as in Example 4.4, so $\{\widehat{\beta^{2k+1}}\}$ is an infinite family of prime hyperbolic knots as in Theorem 1. However, $\widehat{\beta^{2k+1}}$ are not highly twisted plats, as so we cannot explicitly calculate the Hempel distance.